OpenStudy (nottim):

Solve the separable differential equation. y' = 3 y^2

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}=3y^2\\\frac{1}{y^2}dy=3dx\\\int\frac{1}{y^2}dy=\int3dx\\-\frac{1}{y}=3x+c\\y=\frac{-1}{3x+c}\)

OpenStudy (zzr0ck3r):

make sense?

OpenStudy (nottim):

my answer looks nothing likethat...

OpenStudy (nottim):

Oh wait, forgot this: Using the initial condition: y(3)= 3, find y(1). y(1) =

OpenStudy (nottim):

but even then, i still don't understand

OpenStudy (zzr0ck3r):

\(y(3)=3\implies 3=\frac{-1}{3*3+c}\implies27+3c=-1\implies3c=-28 \) so \(c=\frac{-28}{3}\) so \(y(x)=\frac{-1}{3x+\frac{28}{3}}=\frac{-1}{\frac{9x+28}{3}}=\frac{-3}{9x+28}\) now plug in x=1

OpenStudy (zzr0ck3r):

what part of the first part do you not get?

OpenStudy (nottim):

Like, what you did follows what I see in the textbook, but I stil can't follow it.

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}=3y^2\\\frac{1}{y^2}dy=3dx\) are you good up till here?

OpenStudy (nottim):

Ok

OpenStudy (zzr0ck3r):

then we take the integral \(\int\frac{1}{y^2}dy=\int3dx\)

OpenStudy (nottim):

\[\int\limits_{?}^{?}dyy ^{-1}+c=\int\limits_{?}^{?}3xdx+c\]

OpenStudy (zzr0ck3r):

?

OpenStudy (zzr0ck3r):

where did y^(-1) come from, and the + c will come after you take the integral

OpenStudy (nottim):

y^-1 is equiv to 1/y right?

OpenStudy (zzr0ck3r):

yes but we are looking at 1/y^2

OpenStudy (zzr0ck3r):

ok i see what you did.

OpenStudy (zzr0ck3r):

your integral sign shuold be gone, you have taken the intgral, also the dy and dx

OpenStudy (nottim):

why is it -1/y instead of 1/y for your first ever response answer?

OpenStudy (zzr0ck3r):

you should have y^(-1) + c = 3x+c, and technically these are difference c's but we can just add them together and call them c so y^(-1) + c = 3x

OpenStudy (zzr0ck3r):

go the other way and you will see take the derivative of -y^(-1)

OpenStudy (nottim):

derive, not integral?

OpenStudy (zzr0ck3r):

if you want to see where that negative came from...

OpenStudy (zzr0ck3r):

\(\int x^n dx=\frac{x^{n+1}}{n+1}+c\)

OpenStudy (nottim):

i got y^-2

OpenStudy (zzr0ck3r):

exactly

OpenStudy (zzr0ck3r):

thats why the negative is there....

OpenStudy (zzr0ck3r):

so for you y^(-2) becomes y^(-1) / (-1) = -y^(-1)

OpenStudy (nottim):

I see